### v1

# A Methodological Report on Anomaly Detection on Dynamic Knowledge Graphs

Abstract
In this paper, we explore different approaches to anomaly detection on dynamic knowledge graphs, specifically in a microservices environment for Kubernetes applications. Our approach explores three dynamic knowledge graph representations: sequential data, one-hop graph structure, and two-hop graph structure, with each representation incorporating increasingly complex structural information. Each phase includes different machine learning and deep learning models. We empirically analyse their performance and propose an approach based on ensemble learning of these models. Our approach significantly outperforms the baseline on the ISWC 2024 Dynamic Knowledge Graph Anomaly Detection dataset, providing a robust solution for anomaly detection in dynamic complex data.
Keywords:
Dynamic Knowledge Graph Anomaly Detection.
1 Introduction
With the development of graph machine learning and geometric deep learning [1], research on encoding and prediction has been applied to more domains than ever before. Graphs can simultaneously represent entities through nodes [^1] Also referred to as ”vertices”. and relationships between entities through edges. As a result, graph representations are widely used in domains as diverse as transportation [2], chemistry [3], and social networks [24]. Two areas of relative difficulty in graph learning are heterogeneous graphs and dynamic graphs. Heterogeneous graphs allow for different types of nodes and/or edges in the graph [4]. They can therefore better model complex systems with different types of entities, such as a distributed information system with services, pods and nodes. Dynamic graphs allow the structure and/or attributes of a graph to change over time [5], allowing better modelling of data with both temporal and structural information.
Dynamic Knowledge Graphs (DKGs), as an intersection of heterogeneous and dynamic graphs, have attracted increasing attention due to their ability to represent the evolving nature of knowledge over time. By exploiting the ability to capture topological and attribute transformations, DKGs provide valuable insights into the evolution of knowledge in diverse domains such as social media [23], the Internet of Things (IoT) [22, 21], and e-commerce [^2] https://eval.ai/web/challenges/challenge-page/2267/overview.
On the leading Microservices Architecture (MSA) platform, Kubernetes, there are different types of entities such as compute nodes, deployable units (pods), and services that work together to build applications such as e-commerce platforms. These data are typically stored in DKGs in the TTL format. Predictive analysis based on such data is critical to the operation and optimisation of these platforms, particularly for anomaly detection tasks aimed at accurately identifying various types of errors and anomalies.
Despite rapid progress in commercial applications, as a relatively young field, there is still a lack of academic datasets and methodologies. This scarcity can be attributed to several challenges: anomaly events are typically very rare, resulting in an unbalanced data distribution that makes it difficult for models to effectively learn anomaly features and affects their generalisation ability in detecting new anomalies. In addition, models need to capture temporal and structural information and handle heterogeneous nodes/edges, which makes common RNNs and homogeneous GNNs inadequate.
To tackle these challenges, this paper proposes a new approach to anomaly detection on dynamic knowledge graphs using graph structure and temporal dependencies. This approach combines neighbourhood feature extraction with support vector machines (SVMs), extreme gradient boosting (XGB), and self-attention mechanisms. Our approach significantly outperforms the baseline performance on the ’Dynamic Knowledge Graph Anomaly Detection’ challenge dataset at ISWC 2024, providing a robust solution for anomaly detection in dynamic, complex environments.
The structure of the paper is as follows. Section 2 gives the formulation of the problem and describes related work. Section 3 presents the methodology of the paper, while section 4 describes the experimental part of the work and discusses the results obtained. Section 5 concludes our contribution to the ADDKG challenge and discusses future work.
2 Related Work
Recent studies have shown that the dynamic knowledge graph representation is expressive for anomaly detection in temporally complex systems. In this section, we review the related work in this area. In subsection 2.1, we present the definition and notations of the Dynamic Knowledge Graph. Then, in subsection 2.2, we present common approaches to anomaly detection.
2.1 Dynamic Knowledge Graphs
In static graphs, common representations include edge lists and adjacency matrices. Among them, the edge list is widely used in software development as it requires less memory. An edge list is a collection of edges, each represented as a pair of nodes.
In the realm of dynamic graphs, existing research categorises them into Continuous-Time Dynamic Graphs (CTDG) and Discrete-Time Dynamic Graphs (DTDG). CTDGs represent dynamic graphs using timestamped events, such as $(u,v,t)$, where $u$ and $v$ are nodes and $t$ is the timestamp of the interaction. In contrast, DTDGs represent dynamic graphs as a sequence of static graph snapshots $\{G_{1},G_{2},...,G_{T}\}$ at different time intervals.
Combined with the dynamic graph representations, the DKG is a heterogeneous dynamic graph represented by timestamped edge lists. Among the DKG representation formats, TTL (Turtle) is commonly used. A typical TTL allows for the serialisation of triples, which can include nodes, edges, and their attributes, along with timestamps, making TTL versatile for representing entities and relations in an MSA system. An example of a TTL-based DKG is given in Eq.1, where $s$ is the subject node, $o$ is the object node, $p$ is the predicate (or relationship), and $t$ is the timestamp.
$\text{DKG}=\{(s,p,o,t)\mid s,o\in\text{Nodes},p\in\text{Predicate},t\in\text{Timestamp}\}$ (1)
In the problem studied in this paper, anomalous data is represented by its entity $s$, the time interval during which the anomaly occurs $(t_{start},t_{end})$, and the class of the anomaly $c$.
2.2 Anomaly Detection
Anomaly detection is a crucial task in various fields, including network security, fraud detection, and financial systems [25, 26]. This task focuses on identifying anomalous patterns that deviate from the typical observation of the data [15].
Traditional anomaly detection methods typically rely on statistical techniques and distance-based measures. However, these methods often struggle with high-dimensional data and complex dependencies [27]. To overcome these problems, common machine learning approaches include support vector machines (SVM) [6], isolation forests (IF) [7], and extreme gradient boosting (XGB) [8]. SVM [6] identifies the optimal hyperplane that maximises the distance between normal data points and the decision boundary, effectively distinguishing outliers. Isolation Forest [7] is an unsupervised method and detects anomalies by constructing decision trees to isolate observations, with outliers requiring fewer splits to isolate due to their rarity and distinctiveness. XGB [8] is a gradient boosting decision tree algorithm that is able to evaluate the importance of each feature and find those that are most effective in distinguishing normal from abnormal data.
With the advancement of software algorithms and hardware devices, deep learning methods have been increasingly applied to anomaly detection due to their ability to extract complex features [16]. Common architectures for time series include recurrent neural networks (RNN) [9], temporal convolutional networks (TCN) [12], and the self-attention mechanism (SA) [13]. RNN (Recurrent Neural Network) is a type of neural network designed to process sequential data by capturing temporal dependencies. LSTM (Long Short-Term Memory) [10] and GRU (Gated Recurrent Unit) [11] are variants of RNNs designed to solve the vanishing and exploding gradient problems in long sequence training. TCN (Temporal Convolutional Network) [12] captures long-term temporal dependencies with causal convolutions. Self-attention mechanisms capture long-range dependencies by computing the relationships between all elements in a sequence, making them particularly effective for identifying complex patterns. For graph anomaly detection, Graph Neural Networks (GNNs) propagate information through graph structures to identify anomalous patterns [19, 18]. Despite the aforementioned studies proposing numerous algorithms for anomaly detection, significant challenges remain when applied to dynamic knowledge graphs.
2.3 Challenges of Anomaly Detection on Dynamic Knowledge Graph
Due to the complex temporal and structural dependencies in dynamic knowledge graphs, traditional machine learning methods struggle to capture the underlying features. However, deep learning models typically require large amounts of labelled data for effective training, and anomalous data is inherently scarce.
Although anomaly detection tasks on dynamic graphs without attributed edges can generate anomalous edges through methods such as negative sampling [19, 18, 20], this approach becomes infeasible in the case of dynamic knowledge graphs with attributed edges. Therefore, effectively capturing the temporal and structural dependencies of features on dynamic knowledge graphs is crucial to tackling this challenge.
3 Methodology
This section first introduces three representations for dynamic knowledge graphs studied in this task: sequential data, one-hop graph structure, and two-hop graph structure, with each representation incorporating increasingly complex structural information. We then describe the machine learning/deep learning methods applied to each representation. At the end of the section, we propose ensemble schemes for combining multiple methods.
3.1 DKG as Sequential Data
The time-evolving attributes of nodes can be obtained by extracting raw TTL data. For example, in MSA, common node categories could be $\mathcal{C}$ representing "Cluster", $\mathcal{P}$ representing "Pod", $\mathcal{S}$ representing "Service", and $\mathcal{N}$ representing "Node". For each given node category $k$, the time-evolving attributes can be represented as a tensor $\mathbf{\tilde{X}}_{k}\in\mathbb{R}^{N_{k}\times T\times d_{k}}$, where $N_{k}$ indicated the number of instances of the category $k$, $T$ indicates the number of snapshots, and $d_{k}$ indicates the number of features.
$\mathbf{f}(t_{j})=[\mathbf{x}(t_{j}-\tau+1)\,|\,\mathbf{x}(t_{j}-\tau+2)\,|\,\cdots\,|\,\mathbf{x}(t_{j})]$ (2)
Therefore, the sequential data representation of a given category in the DKG can be expressed by a matrix $\mathbf{F}\in\mathbb{R}^{T\times D_{k}}$ as shown in Eq. 3, where $\mathbf{f}(t_{j})$ is the aggregated attribute vector at timestamp $t_{j}$.
$\mathbf{F}=\begin{bmatrix}\mathbf{f}(t_{1})\\ \mathbf{f}(t_{2})\\ \vdots\\ \mathbf{f}(t_{T})\end{bmatrix}=\begin{bmatrix}\mathbf{x}(t_{1}-\tau+1)&\mathbf{x}(t_{1}-\tau+2)&\cdots&\mathbf{x}(t_{1})\\ \mathbf{x}(t_{2}-\tau+1)&\mathbf{x}(t_{2}-\tau+2)&\cdots&\mathbf{x}(t_{2})\\ \vdots&\vdots&\ddots&\vdots\\ \mathbf{x}(t_{T}-\tau+1)&\mathbf{x}(t_{m}-\tau+2)&\cdots&\mathbf{x}(t_{m})\end{bmatrix}$ (3)
Once the matrix representation of the attributes $\mathbf{F}$ is obtained, additional information can be calculated. For example, the difference captures the trend of the time series, while the variance reflects the stability of the data fluctuations. Specific formulas for statistical properties are given in Eq. 4.
$\text{Difference}(t)$ $=$ $Z(t){-}Z(t{-}1)$ (4)
$\text{Variance}(t)$ $=$ $\frac{1}{\tau}\sum_{j=t-\tau+1}^{t}(\mathbf{Z}(j)-\bar{\mathbf{Z}})^{2}$
$\text{Rolling Mean}_{\tau}(t)$ $=$ $\frac{1}{\tau}\sum_{j=t-\tau+1}^{t}\mathbf{Z}(j)$
$\text{Rolling Sum}_{\tau}(t)$ $=$ $\sum_{j=t-\tau+1}^{t}\mathbf{Z}(j)$
Based on this data representation, the temporal information across multiple time steps is aggregated into a single vector. Consequently, machine learning algorithms such as SVM, XGB, and IF, as well as deep learning models like MLP, TCN, and RNN, can use this input to capture information across multiple time steps.
3.2 One-Hop Graph Structure
To explore the structural information of the knowledge graph, we propose to use one-hop neighbour aggregation to capture the relationships between different entities. Given that the relationships in the DKG of an MSA are mainly hierarchical, i.e., a cluster contains multiple nodes, a node contains multiple pods, and a pod is connected to services through multiple connections (see Fig.1), it is possible to aggregate information from leaf nodes to root nodes to obtain the structural-temporal features of each node at each timestamp based on the temporal information considered in the previous subsection.
$S(t)=\bigcup_{i=1}^{n_{S}}\left(S_{i}(t),\bigcup_{j\in C(S_{i})}\left(C_{j}(t),\bigcup_{k\in P(C_{j})}\left(P_{k}(t),\bigcup_{l\in N(P_{k})}N_{l}(t)\right)\right)\right)$ (5)
For instance, to encode and predict data for the category Service, information related to the Service can be aggregated. Specifically, for all service instances $\mathcal{S}(t)$ at timestamp $t$, the dynamic information from related Nodes, Pods, and Connections can be aggregated to form the attributes of the Service, as shown in Eq. 5 where $S_{i}$ represents the $i$-th instance of the Service category, $C_{j}(S_{i})$ represents the set of Connection instances related to $S_{i}$, $P_{k}(C_{j})$ represents the set of Pod instances related to $C_{j}$, and $N_{l}(P_{k})$ represents the set of Node instances related to $P_{k}$. In this manner, multiple one-hop graph structures are used for message propagation, and these structural-temporal features provide the model with a richer context.
Figure 1: The dataset consists of knowledge graphs representing an e-commerce Kubernetes cluster at specific timestamps (UTC). Each snapshot provides information about the state of the microservice platform, deployments, resources, and component relationships. The graphs are in TTL format and detail subjects, objects, and their relationships based on a defined ontology. The ontology overview illustrates the relationships between Kubernetes concepts.
3.3 Two-Hop Graph Structure
Some graph machine learning algorithms have pointed out that a larger convolutional field, or aggregation of multi-hop neighbours, can optimise the model. Inspired by this, we realised that aggregating second-hop neighbours could provide richer information. In the tree-like graph of the MSA problem, considering only the current level and its associated root level forms a bipartite graph. For nodes of category $k$, the two-hop neighbours on this bipartite graph are still nodes of the same category.
Recognising that there may be interdependencies between Service instances at each timestamp, we propose to use two-hop neighbour aggregation. Intuitively, this aggregation can be achieved by concatenating vertex features, as shown in Eq. 6 where $N_{\mathcal{S}}$ represents the number of Service instances and $S_{i}(t)$ represents the data for the $i$-th Service instance at timestamp $t$.
$\mathbf{X}_{\text{service}}(t)=[\mathbf{S_{1}}(t),\mathbf{S_{2}}(t),\ldots,\mathbf{S_{N_{S}}}(t)];$ (6)
Another method of capturing two-hop neighbours is using the self-attention mechanism to capture the correlations among vertices. The self-attention mechanism works by computing the relevance (or attention) between each pair of elements in a sequence, enabling the model to weigh the importance of each element relative to others. This allows the model to focus on relevant parts of the input sequence when making predictions, effectively capturing long-range dependencies and complex relationships within the data.
3.4 Ensemble Learning
Ensemble learning is a prevalent methodology in the domain of machine learning that mitigates variance through the consolidation of multiple models, thereby enhancing model robustness. The integration of diverse and superior-performing models is a prerequisite for the efficacy of ensemble learning [14]. Ensemble learning allows multiple models to make predictions on the same problem and thus vote on the final prediction. The voting principle can be hard voting or soft voting, while the voting mechanism usually consists of unanimous voting and majority voting, which will be introduced in this subsection.
In hard voting, each classifier votes for a class label, and the final prediction is determined by the majority of these classifiers. In soft voting, each classifier provides a probability for each class, and the final prediction is based on the highest average probability across all classifiers.
As two common voting mechanisms combined with hard voting, especially in anomaly detection, unanimous voting classifies an instance as an anomaly only if all the models agree. Majority voting, on the other hand, classifies an instance as an anomaly if the majority of models agree.
Intuitively, unanimous voting is more conservative and reduces false positives for anomaly detection, but may also miss some true anomalies. The majority voting method, in contrast, has some flexibility and can improve the detection of true anomalies, but it also preserves the false alarms. Both methods are tested on multiple sets of models and it are discussed in the next section.
4 Experiment
4.1 Experimental Setup
All experiments were conducted utilising the ISWC 2024 ADDKG dataset. The dataset contains five days of labelled data, with three days of data used for training and the remaining two days used for testing. The data was collected at approximately 15 second intervals and presents a challenge due to its imbalance, with anomalies representing only approximately 4% of the total data. To mitigate this imbalance, class weights were applied to the loss functions during model training to ensure that the models effectively learned to detect anomalies.
As outlined in the methodology section, our approach included three levels of representation. For the first representation, sequential modelling is employed. The original training set is randomly divided into two subsets: 80% for training, and the other 20% for validation. For the other two representations, a more sophisticated approach to data splitting is under consideration. For normal class, the 80-20 split mechanism remains applicable. However, in the case of the anomaly class, it is necessary to identify complete and continuous sequences of anomalous records as "anomalous events". The objective was to maintain the anomalous events in their entirety during the process of data splitting, the splitting of a single anomalous process between the training and validation sets. This approach ensures the contextual integrity of the anomalies, allowing the models to learn more effectively from complete anomalous events. To achieve this, for each anomaly type, one complete event was randomly selected for validation, while the remaining events were used for training.
In the paper, we focus on binary anomaly detection, aiming to distinguish between normal and anomalous data without delving into the specific types of anomalies. The predictions are evaluated by F1 score, precision, and recall.
4.2 Performance of Models
The experimental results of each representation are shown in tables 1 and 2. Table 1 focuses on the performance of individual models, while table 2 highlights the performance of different ensemble methods. D1, D2, and D3 represent sequential data representation, one-hop graph representation, and two-hop graph representation, respectively.
Table 1 shows how each model performs under different structural conditions of the dataset. The results show that the inclusion of graph structures and service interactions generally improves model performance, particularly in terms of recall and F1 score.
Table 1: Performance metrics for each model across different phases. The best performances are highlighted in bold.
Model Dataset Phase F1 Score Precision Recall
Baseline - 0.06759 0.03498 1.00000
MLP D1 0.09390 0.05102 0.58824
XGB D2 0.21622 0.20000 0.23529
SVM D2 0.16185 0.08974 0.82353
TCN D1 0.09816 0.05479 0.47059
D2 0.07792 0.04206 0.52941
LSTM D1 0.06759 0.03498 1.00000
D2 0.10046 0.05446 0.64706
D3 0.16923 0.09735 0.64706
GRU D1 0.06759 0.03498 1.00000
D2 0.07362 0.04110 0.35294
D3 0.16250 0.09091 0.76471
Self-Attention D1 0.07203 0.03736 1.00000
D2 0.09028 0.04797 0.76471
D3 0.22018 0.13043 0.70588
In order to effectively combine multiple models, the second part of the experiments evaluates the performance of different ensemble methods, as shown in Table 2. Different ensemble strategies are tested, such as combining XGB, SVM, stacked layers of the Self Attention model, and Isolation Forest. In the evaluation of the ensemble methods, we have compared two different methods, soft voting, hard&unanimous voting and hard&majority voting. Table 2 shows the best performances obtained. Overall, the soft voting approach provided better results for our objectives.
Table 2: Performance metrics for ensemble methods across different phases.The best performance are highlighted in bold.
Ensemble Model Method Dataset Phase F1 Score Precision Recall
Baseline - - 0.06759 0.03498 1.00000
XGB + IF soft D2 0.34286 0.33333 0.35294
SVM + IF hard,unanimous D2 0.19802 0.11905 0.58824
XGB + SVM + IF soft D2 0.38596 0.275 0.64706
XGB + SVM + SA + IF hard,unanimous D2+D3 0.36364 0.2963 0.47059
XGB + SVM + SA soft D2+D3 0.51429 0.5 0.52941
The individual models using the three representations mentioned in Section 3 and the results using ensemble learning both significantly outperforms the baseline result  [^3] The baseline result is from https://eval.ai/web/challenges/challenge-page/2267/overview, demonstrating the potential of our approach.
5 Conclusion
In this paper, we investigated the effectiveness of different machine learning and deep learning models on the challenging task of anomaly detection on dynamic knowledge graphs, with a specific focus on a microservices environment. Using a structured experimental approach, we analysed the performance of the models in three different representations: sequential data modelling, one-hop graph structure, and two-hop graph structure.
Our results show that incorporating graph structures significantly improves the performance of anomaly detection models. In particular, the Two-Hop Graph Structure, which includes service interactions, produced the best results across most metrics, highlighting the importance of capturing complex interdependencies within the data. Deep learning models, particularly the Two-Stacked Self-Attention (2StackedSA) model, showed superior performance in leveraging these structures.
Furthermore, the ensemble methods provided a substantial boost in performance by combining the strengths of top-performing individual models. Among the ensemble strategies, the Unanimous Voting approach, which requires all models to agree on the prediction of an anomaly, outperformed the Majority Voting approach in terms of precision. This aligns well with the competition’s emphasis on precision, highlighting the effectiveness of this conservative yet accurate method.
Future work should address several areas for improvement. Firstly, improving the scalability of our models to handle larger and more complex datasets would be crucial. Secondly, exploring more advanced ensemble techniques, such as stacking or blending, could further improve detection accuracy. Thirdly, the integration of real-time anomaly detection capabilities would be highly beneficial for practical applications in dynamic environments. Finally, our study was limited to binary classification (normal/abnormal). Future research should focus on improving the accuracy of predicting different types of anomalies. Furthermore, although we used graph structure information, we did not apply Graph Neural Networks (GNNs), which could be a promising direction for future investigation and testing.
This study demonstrates the significant potential of combining graph-based data representations with advanced machine learning and deep learning techniques for anomaly detection on dynamic knowledge graphs. By refining these methods and exploring the future directions outlined, we can develop more robust, accurate and scalable solutions for complex anomaly detection tasks in various domains.

### v2

@@
 Dynamic graphs allow the structure and/or attributes of a graph to change over time [5], allowing better modelling of data with both temporal and structural information.
-Dynamic Knowledge Graphs (DKGs), as an intersection of heterogeneous and dynamic graphs, have attracted increasing attention due to their ability to represent the evolving nature of knowledge over time.
+Dynamic Knowledge Graphs (DKG), as an intersection of heterogeneous and dynamic graphs, have attracted increasing attention due to their ability to represent the evolving nature of knowledge over time.
 By exploiting the ability to capture topological and attribute transformations, DKGs provide valuable insights into the evolution of knowledge in diverse domains such as social media [23], the Internet of Things (IoT) [22, 21], and e-commerce [^2] https://eval.ai/web/challenges/challenge-page/2267/overview.
-On the leading Microservices Architecture (MSA) platform, Kubernetes, there are different types of entities such as compute nodes, deployable units (pods), and services that work together to build applications such as e-commerce platforms.
-These data are typically stored in DKGs in the TTL format.
-Predictive analysis based on such data is critical to the operation and optimisation of these platforms, particularly for anomaly detection tasks aimed at accurately identifying various types of errors and anomalies.
+On the leading Micro-Services Architecture (MSA) platform, such as Kubernetes, there are different types of entities such as compute nodes, deployable units (pods), and services that work together to build applications such as e-commerce platforms.
+These data are typically stored in DKGs in the TTL format [^3] https://en.wikipedia.org/wiki/Turtle_(syntax).
+Predictive analysis based on such data is critical to the operation and optimisation of these platforms, particularly for anomaly detection tasks aimed at accurately identifying various types of anomalies.
 Despite rapid progress in commercial applications, as a relatively young field, there is still a lack of academic datasets and methodologies.
@@
 In addition, models need to capture temporal and structural information and handle heterogeneous nodes/edges, which makes common RNNs and homogeneous GNNs inadequate.
-To tackle these challenges, this paper proposes a new approach to anomaly detection on dynamic knowledge graphs using graph structure and temporal dependencies.
-This approach combines neighbourhood feature extraction with support vector machines (SVMs), extreme gradient boosting (XGB), and self-attention mechanisms.
-Our approach significantly outperforms the baseline performance on the ’Dynamic Knowledge Graph Anomaly Detection’ challenge dataset at ISWC 2024, providing a robust solution for anomaly detection in dynamic, complex environments.
+To address these challenges, this paper introduces a novel anomaly detection framework for dynamic knowledge graphs, leveraging both graph structure and temporal dependencies.
+Our method integrates neighbourhood feature extraction with advanced machine learning techniques, including Support Vector Machine (SVM) [6], Extreme Gradient Boosting (XGB) [8], and Self-Attention mechanism (SA) [13] to enhance the model’s ability to capture complex patterns of anomalies.
+Our approach achieves significant performance improvements over the baseline in the Anomaly Detection on Dynamic Knowledge Graphs challenge (ADDKG) [^4] https://eval.ai/web/challenges/challenge-page/2267/overview at ISWC 2024, providing a robust solution for anomaly detection in dynamic and complex environments.
 The structure of the paper is as follows.
 Section 2 gives the formulation of the problem and describes related work.
-Section 3 presents the methodology of the paper, while section 4 describes the experimental part of the work and discusses the results obtained.
+Section 3 presents the methodology of the paper.
+Section 4 describes the experimental part of the work and discusses the results obtained.
 Section 5 concludes our contribution to the ADDKG challenge and discusses future work.
@@
 In contrast, DTDGs represent dynamic graphs as a sequence of static graph snapshots $\{G_{1},G_{2},...,G_{T}\}$ at different time intervals.
-Combined with the dynamic graph representations, the DKG is a heterogeneous dynamic graph represented by timestamped edge lists.
-Among the DKG representation formats, TTL (Turtle) is commonly used.
-A typical TTL allows for the serialisation of triples, which can include nodes, edges, and their attributes, along with timestamps, making TTL versatile for representing entities and relations in an MSA system.
-An example of a TTL-based DKG is given in Eq.1, where $s$ is the subject node, $o$ is the object node, $p$ is the predicate (or relationship), and $t$ is the timestamp.
-$\text{DKG}=\{(s,p,o,t)\mid s,o\in\text{Nodes},p\in\text{Predicate},t\in\text{Timestamp}\}$ (1)
+Combined with the dynamic graph representations, the Dynamic Knowledge Graph (DKG) is modelled as a heterogeneous dynamic graph, represented by timestamped edge lists.
+Among the various formats for representing DKGs, Turtle (TTL) is one of the most commonly used.
+TTL allows the serialisation of nodes, edges, their attributes and timestamps, making it a versatile format for representing entities and relationships in a multi-agent system (MSA).
+An example of a TTL-based DKG is given in Eq.
+1, where $s$ is the subject node, $o$ is the object node, $p$ is the predicate (or relationship), and $t$ is the timestamp.
+$\text{DKG}=\{(s,p,o,t)\mid t\in[t_{1},t_{2},...,T]\}$ (1)
 In the problem studied in this paper, anomalous data is represented by its entity $s$, the time interval during which the anomaly occurs $(t_{start},t_{end})$, and the class of the anomaly $c$.
+Table 1 shows all the anomaly types $c$ in the dataset.
+Table 1: Description of system faults and their types.
+Fault Type Description
+Node CPU Hog Static Consumes the CPU resources of the targeted node.
+Node Drain Dynamic Forces rescheduling of the resources running on
+the targeted node.
+Node Memory Hog Static Consumes the memory resources of the targeted
+node.
+Pod Autoscaler Dynamic Changes the desired number of replicas for the
+targeted workload.
+Pod CPU Hog Static Consumes the CPU resources of the targeted
+application container.
+Pod Delete Dynamic Causes forced/graceful pod failure of a random
+replica of the targeted workload
+Pod Network Latency Static Latency introduced in the network communication
+between different pods.
 2.2 Anomaly Detection
@@
 However, these methods often struggle with high-dimensional data and complex dependencies [27].
-To overcome these problems, common machine learning approaches include support vector machines (SVM) [6], isolation forests (IF) [7], and extreme gradient boosting (XGB) [8].
-SVM [6] identifies the optimal hyperplane that maximises the distance between normal data points and the decision boundary, effectively distinguishing outliers.
-Isolation Forest [7] is an unsupervised method and detects anomalies by constructing decision trees to isolate observations, with outliers requiring fewer splits to isolate due to their rarity and distinctiveness.
-XGB [8] is a gradient boosting decision tree algorithm that is able to evaluate the importance of each feature and find those that are most effective in distinguishing normal from abnormal data.
-With the advancement of software algorithms and hardware devices, deep learning methods have been increasingly applied to anomaly detection due to their ability to extract complex features [16].
+To overcome these problems, common machine learning approaches include Support Vector Machine (SVM) [6], eXtreme Gradient Boosting (XGB) [8], and Isolation Forest (IF) [7].
+SVM identifies the optimal hyperplane that maximises the distance between normal data points and the decision boundary, effectively distinguishing outliers.
+XGB is a gradient-boosting decision tree algorithm that is able to evaluate the importance of each feature and find those that are most effective in separating normal from abnormal data.
+Isolation Forest is an unsupervised method that detects anomalies by constructing decision trees to isolate observations, with outliers requiring fewer splits to isolate due to their rarity and distinctiveness.
+With the advancement of software algorithms and hardware, deep learning methods have been increasingly popular for anomaly detection due to their ability to extract complex features [16].
 Common architectures for time series include Recurrent Neural Networks (RNN) [9], Temporal Convolutional Networks (TCN) [12], and the Self-Attention mechanism (SA) [13].
-RNN (Recurrent Neural Network) is a type of neural network designed to process sequential data by capturing temporal dependencies.
-LSTM (Long Short-Term Memory) [10] and GRU (Gated Recurrent Unit) [11] are variants of RNNs designed to solve the vanishing and exploding gradient problems in long sequence training.
-TCN (Temporal Convolutional Network) [12] captures long-term temporal dependencies with causal convolutions.
-Self-attention mechanisms capture long-range dependencies by computing the relationships between all elements in a sequence, making them particularly effective for identifying complex patterns.
-For graph anomaly detection, Graph Neural Networks (GNNs) propagate information through graph structures to identify anomalous patterns [19, 18].
-Despite the aforementioned studies proposing numerous algorithms for anomaly detection, significant challenges remain when applied to dynamic knowledge graphs.
-2.3 Challenges of Anomaly Detection on Dynamic Knowledge Graph
-Due to the complex temporal and structural dependencies in dynamic knowledge graphs, traditional machine learning methods struggle to capture the underlying features.
-However, deep learning models typically require large amounts of labelled data for effective training, and anomalous data is inherently scarce.
-Although anomaly detection tasks on dynamic graphs without attributed edges can generate anomalous edges through methods such as negative sampling [19, 18, 20], this approach becomes infeasible in the case of dynamic knowledge graphs with attributed edges.
-Therefore, effectively capturing the temporal and structural dependencies of features on dynamic knowledge graphs is crucial to tackling this challenge.
+RNNs, including Long Short-Term Memory (LSTM) [10] and Gated Recurrent Unit (GRU) [11], capture temporal dependencies, addressing gradient issues in long sequences.
+TCNs handle long-term dependencies through causal convolutions, while SA captures long-range dependencies by examining relationships across sequences.
+For graph anomaly detection, Graph Neural Networks (GNN) propagate information through graph structures to identify anomalous patterns [19, 18].
+Despite advancements in anomaly detection algorithms, significant challenges persist in dynamic knowledge graphs.
+Traditional machine learning methods struggle to capture the complex temporal and structural dependencies, while deep learning models require large amounts of labelled data, which is often scarce.
+Although techniques like negative sampling [19, 18, 20] can generate anomalous edges in dynamic graphs without attributed edges, this approach is infeasible for dynamic knowledge graphs with attributed edges.
+Capturing both temporal and structural dependencies remains key to tackling this challenge.
 3 Methodology
-This section first introduces three representations for dynamic knowledge graphs studied in this task: sequential data, one-hop graph structure, and two-hop graph structure, with each representation incorporating increasingly complex structural information.
-We then describe the machine learning/deep learning methods applied to each representation.
-At the end of the section, we propose ensemble schemes for combining multiple methods.
+This section introduces three representations of dynamic knowledge graphs: sequential data, one-hop, and two-hop graph structures, each adding increasing structural complexity.
+We then describe the machine learning and deep learning methods applied to these representations, followed by proposed ensemble learning schemes.
+Figures 1 and 2 provide an overview of the pipeline, including the data structure, the three levels of the data processing, and the final output.
+Figure 1: Left: The ontology overview of the relationships between Kubernetes concepts of the ADDKG dataset.
+Right: The simplified hierarchical structure of the ADDKG dataset.
+The container information is aggregated as the attributes of Pods.
+The Connections serve as attributes of the edges between Pods and Services.
+Figure 2: This figure presents the methodology used in this study, with three levels of processing: sequential data, one-hop, and two-hop graph structures.
+Each level introduces increasing structural complexity, showing how these representations are fed into anomaly detection models.
 3.1 DKG as Sequential Data
-The time-evolving attributes of nodes can be obtained by extracting raw TTL data.
-For example, in MSA, common node categories could be $\mathcal{C}$ representing "Cluster", $\mathcal{P}$ representing "Pod", $\mathcal{S}$ representing "Service", and $\mathcal{N}$ representing "Node".
+The time-evolving attributes of nodes can be obtained by extracting raw TTL data, without considering the structural information of the DKG.
+As shown in the "DKG as Sequential Data" part of Fig.
+2, we merge all the sequential data of all the vertices in a DKG by the timestamp.
+This includes vertices of type $\mathcal{C}$ representing "Cluster", type $\mathcal{P}$ representing "Pod", type $\mathcal{S}$ representing "Service", and type $\mathcal{N}$ representing "Node".
 For each given node category $k$, the time-evolving attributes can be represented as a tensor $\mathbf{\tilde{X}}_{k}\in\mathbb{R}^{N_{k}\times T\times d_{k}}$, where $N_{k}$ indicated the number of instances of the category $k$, $T$ indicates the number of snapshots, and $d_{k}$ indicates the number of features.
+A sliding window then captures temporal dependencies by concatenating the historical information of the past $\tau$ time steps, as shown in Eq.
+2.
 $\mathbf{f}(t_{j})=[\mathbf{x}(t_{j}-\tau+1)\,|\,\mathbf{x}(t_{j}-\tau+2)\,|\,\cdots\,|\,\mathbf{x}(t_{j})]$ (2)
-Therefore, the sequential data representation of a given category in the DKG can be expressed by a matrix $\mathbf{F}\in\mathbb{R}^{T\times D_{k}}$ as shown in Eq.
+Therefore, the sequential data representation of a given category $k$ in the DKG can be expressed by a matrix $\mathbf{X}\in\mathbb{R}^{T\times D_{k}}$, where $D_{k}=N_{k}\times d_{k}$, as shown in Eq.
 3, where $\mathbf{f}(t_{j})$ is the aggregated attribute vector at timestamp $t_{j}$.
-$\mathbf{F}=\begin{bmatrix}\mathbf{f}(t_{1})\\ \mathbf{f}(t_{2})\\ \vdots\\ \mathbf{f}(t_{T})\end{bmatrix}=\begin{bmatrix}\mathbf{x}(t_{1}-\tau+1)&\mathbf{x}(t_{1}-\tau+2)&\cdots&\mathbf{x}(t_{1})\\ \mathbf{x}(t_{2}-\tau+1)&\mathbf{x}(t_{2}-\tau+2)&\cdots&\mathbf{x}(t_{2})\\ \vdots&\vdots&\ddots&\vdots\\ \mathbf{x}(t_{T}-\tau+1)&\mathbf{x}(t_{m}-\tau+2)&\cdots&\mathbf{x}(t_{m})\end{bmatrix}$ (3)
-Once the matrix representation of the attributes $\mathbf{F}$ is obtained, additional information can be calculated.
+$\mathbf{X}=\begin{bmatrix}\mathbf{f}(t_{1})\\ \mathbf{f}(t_{2})\\ \vdots\\ \mathbf{f}(t_{T})\end{bmatrix}=\begin{bmatrix}\mathbf{x}(t_{1}-\tau+1)&\mathbf{x}(t_{1}-\tau+2)&\cdots&\mathbf{x}(t_{1})\\ \mathbf{x}(t_{2}-\tau+1)&\mathbf{x}(t_{2}-\tau+2)&\cdots&\mathbf{x}(t_{2})\\ \vdots&\vdots&\ddots&\vdots\\ \mathbf{x}(t_{T}-\tau+1)&\mathbf{x}(t_{m}-\tau+2)&\cdots&\mathbf{x}(t_{m})\end{bmatrix}$ (3)
+Once the matrix representation of the attributes $\mathbf{X}$ is obtained, additional information can be computed.
 For example, the difference captures the trend of the time series, while the variance reflects the stability of the data fluctuations.
-Specific formulas for statistical properties are given in Eq.
-4.
-$\text{Difference}(t)$ $=$ $Z(t){-}Z(t{-}1)$ (4)
-$\text{Variance}(t)$ $=$ $\frac{1}{\tau}\sum_{j=t-\tau+1}^{t}(\mathbf{Z}(j)-\bar{\mathbf{Z}})^{2}$
-$\text{Rolling Mean}_{\tau}(t)$ $=$ $\frac{1}{\tau}\sum_{j=t-\tau+1}^{t}\mathbf{Z}(j)$
-$\text{Rolling Sum}_{\tau}(t)$ $=$ $\sum_{j=t-\tau+1}^{t}\mathbf{Z}(j)$
 Based on this data representation, the temporal information across multiple time steps is aggregated into a single vector.
@@
 To explore the structural information of the knowledge graph, we propose to use one-hop neighbour aggregation to capture the relationships between different entities.
-Given that the relationships in the DKG of an MSA are mainly hierarchical, i.e., a cluster contains multiple nodes, a node contains multiple pods, and a pod is connected to services through multiple connections (see Fig.1), it is possible to aggregate information from leaf nodes to root nodes to obtain the structural-temporal features of each node at each timestamp based on the temporal information considered in the previous subsection.
+Given that the relationships in the DKG of an MSA are mainly hierarchical, i.e., a cluster contains multiple nodes, a node contains multiple pods, and a pod is connected to services through multiple connections (see Fig.
+1), it is possible to aggregate information from leaf nodes to root nodes to obtain the structural-temporal features of each node at each timestamp, as shown the aggregate method in the "One-Hop Chart Structure" section of Fig.
+2.
 $S(t)=\bigcup_{i=1}^{n_{S}}\left(S_{i}(t),\bigcup_{j\in C(S_{i})}\left(C_{j}(t),\bigcup_{k\in P(C_{j})}\left(P_{k}(t),\bigcup_{l\in N(P_{k})}N_{l}(t)\right)\right)\right)$ (4)
 For instance, to encode and predict data for the category Service, information related to the Service can be aggregated.
-Specifically, for all service instances $\mathcal{S}(t)$ at timestamp $t$, the dynamic information from related Nodes, Pods, and Connections can be aggregated to form the attributes of the Service, as shown in Eq.
-5 where $S_{i}$ represents the $i$-th instance of the Service category, $C_{j}(S_{i})$ represents the set of Connection instances related to $S_{i}$, $P_{k}(C_{j})$ represents the set of Pod instances related to $C_{j}$, and $N_{l}(P_{k})$ represents the set of Node instances related to $P_{k}$.
+Specifically, for all service instances $\mathcal{S}(t)$ at timestamp $t$, the dynamic information from related Nodes, Pods, Containers and Connections can be aggregated to form the attributes of the Service, as shown in Eq.
+4, where $n_{S}$ represents the number of Service instances, $S_{i}$ represents the $i$-th instance of the Service category, $C_{j}(S_{i})$ represents the elements in the set of Connection instances related to $S_{i}$, $P_{k}(C_{j})$ represents the elements in the set of Pod instances related to $C_{j}$, and $N_{l}(P_{k})$ represents the elements in the set of Node instances related to $P_{k}$.
 In this manner, multiple one-hop graph structures are used for message propagation, and these structural-temporal features provide the model with a richer context.
-Figure 1: The dataset consists of knowledge graphs representing an e-commerce Kubernetes cluster at specific timestamps (UTC).
-Each snapshot provides information about the state of the microservice platform, deployments, resources, and component relationships.
-The graphs are in TTL format and detail subjects, objects, and their relationships based on a defined ontology.
-The ontology overview illustrates the relationships between Kubernetes concepts.
 3.3 Two-Hop Graph Structure
 Some graph machine learning algorithms have pointed out that a larger convolutional field, or aggregation of multi-hop neighbours, can optimise the model.
-Inspired by this, we realised that aggregating second-hop neighbours could provide richer information.
-In the tree-like graph of the MSA problem, considering only the current level and its associated root level forms a bipartite graph.
-For nodes of category $k$, the two-hop neighbours on this bipartite graph are still nodes of the same category.
-Recognising that there may be interdependencies between Service instances at each timestamp, we propose to use two-hop neighbour aggregation.
-Intuitively, this aggregation can be achieved by concatenating vertex features, as shown in Eq.
+Inspired by the potential of aggregating second-hop neighbours to provide richer information, we observed that in the tree-like graph of the MSA problem, considering only the current level and its root forms a bipartite graph.
+For nodes of category $k$, the two-hop neighbours remain nodes of the same category.
+Recognising potential interdependencies between Service instances at each timestamp, we propose two-hop neighbour aggregation.
+This can be intuitively achieved by concatenating vertex features, as shown in Eq.
 5 where $N_{\mathcal{S}}$ represents the number of Service instances and $S_{i}(t)$ represents the data for the $i$-th Service instance at timestamp $t$.
-$\mathbf{X}_{\text{service}}(t)=[\mathbf{S_{1}}(t),\mathbf{S_{2}}(t),\ldots,\mathbf{S_{N_{S}}}(t)];$ (6)
-Another method of capturing two-hop neighbours is using the self-attention mechanism to capture the correlations among vertices.
-The self-attention mechanism works by computing the relevance (or attention) between each pair of elements in a sequence, enabling the model to weigh the importance of each element relative to others.
+$\mathbf{X}_{\text{service}}(t)=[\mathbf{S_{1}}(t),\mathbf{S_{2}}(t),\ldots,\mathbf{S_{\mathit{N_{S}}}}(t)];$ (5)
+Another method of capturing two-hop neighbours is using the Self-Attention mechanism (SA) to capture the correlations among vertices.
+SA works by computing the relevance (or attention) between each pair of elements in a sequence, enabling the model to weigh the importance of each element relative to others.
 This allows the model to focus on relevant parts of the input sequence when making predictions, effectively capturing long-range dependencies and complex relationships within the data.
+See the "two hop graph structure" part in Fig.
+2.
 3.4 Ensemble Learning
-Ensemble learning is a prevalent methodology in the domain of machine learning that mitigates variance through the consolidation of multiple models, thereby enhancing model robustness.
-The integration of diverse and superior-performing models is a prerequisite for the efficacy of ensemble learning [14].
-Ensemble learning allows multiple models to make predictions on the same problem and thus vote on the final prediction.
-The voting principle can be hard voting or soft voting, while the voting mechanism usually consists of unanimous voting and majority voting, which will be introduced in this subsection.
+Ensemble learning is a prevalent methodology in machine learning that reduce variance and enhance model robustness by combining multiple models.
+The effectiveness of ensemble learning depends on the integration of diverse, high-performing models [14].
+It allows multiple models to make predictions on the same problem and vote on the final predictions.
+Voting can be hard or soft voting, with the mechanism such as unanimous and majority voting, which will be introduced in this subsection.
+Figure 3: The ensemble learning framework used in this competition.
+The best-performing models: XGBoost (XGB), Support Vector Machine (SVM), and Self-Attention (SA) are combined using three voting mechanisms: hard-unanimous, hard-majority, and soft voting, to generate the prediction $\hat{y_{1}}$, where $\hat{y_{1}}\in\{0,1\}$, with 0 indicating normal and 1 indicating anomaly.
+Optionally, a two-stage model combination can be applied, where the unsupervised Isolation Forest further refines the anomalies detected in $\hat{y_{1}}$ to reduce false positives.
 In hard voting, each classifier votes for a class label, and the final prediction is determined by the majority of these classifiers.
 In soft voting, each classifier provides a probability for each class, and the final prediction is based on the highest average probability across all classifiers.
-As two common voting mechanisms combined with hard voting, especially in anomaly detection, unanimous voting classifies an instance as an anomaly only if all the models agree.
+As two common voting mechanisms combined with hard voting, unanimous voting classifies an instance as an anomaly only if all the models agree.
 Majority voting, on the other hand, classifies an instance as an anomaly if the majority of models agree.
@@
 4.1 Experimental Setup
-All experiments were conducted utilising the ISWC 2024 ADDKG dataset.
-The dataset contains five days of labelled data, with three days of data used for training and the remaining two days used for testing.
-The data was collected at approximately 15 second intervals and presents a challenge due to its imbalance, with anomalies representing only approximately 4% of the total data.
-To mitigate this imbalance, class weights were applied to the loss functions during model training to ensure that the models effectively learned to detect anomalies.
-As outlined in the methodology section, our approach included three levels of representation.
-For the first representation, sequential modelling is employed.
-The original training set is randomly divided into two subsets: 80% for training, and the other 20% for validation.
-For the other two representations, a more sophisticated approach to data splitting is under consideration.
-For normal class, the 80-20 split mechanism remains applicable.
-However, in the case of the anomaly class, it is necessary to identify complete and continuous sequences of anomalous records as "anomalous events".
-The objective was to maintain the anomalous events in their entirety during the process of data splitting, the splitting of a single anomalous process between the training and validation sets.
-This approach ensures the contextual integrity of the anomalies, allowing the models to learn more effectively from complete anomalous events.
-To achieve this, for each anomaly type, one complete event was randomly selected for validation, while the remaining events were used for training.
+All experiments were conducted on the ISWC 2024 ADDKG dataset, containing five days of labelled data.
+Three days were used for training and two for testing, with data sampled at 15 second intervals.
+The dataset has a significant imbalance, with anomalies representing only 4% of the total.
+To tackle this, class weights were incorporated into the loss functions to improve the models’ ability of anomaly detection.
+As outlined in the methodology section, we tested three levels of representation.
+For the first representation, sequential modelling is used, where the training set is randomly divided into 80% for training and 20% for validation.
+For the other two representations, a more refined data splitting strategy is used.
+While the 80-20 split is still used for the normal class, anomalies are treated as continuous anomalous events to preserve their integrity.
+This ensures that the models learn from complete sequences of anomalies.
+For each anomaly type, one complete event is randomly selected for validation, and the remaining events are used for training.
 In the paper, we focus on binary anomaly detection, aiming to distinguish between normal and anomalous data without delving into the specific types of anomalies.
@@
 4.2 Performance of Models
-The experimental results of each representation are shown in tables 1 and 2.
-Table 1 focuses on the performance of individual models, while table 2 highlights the performance of different ensemble methods.
-D1, D2, and D3 represent sequential data representation, one-hop graph representation, and two-hop graph representation, respectively.
-Table 1 shows how each model performs under different structural conditions of the dataset.
-The results show that the inclusion of graph structures and service interactions generally improves model performance, particularly in terms of recall and F1 score.
-Table 1: Performance metrics for each model across different phases. The best performances are highlighted in bold.
+This section presents the experimental results for the three data representations: D1, D2, and D3, correspond respectively to sequential data, one-hop, and two-hop graph representations.
+Table 2 summarises the performance of individual models, while table 3 focuses on the results of different ensemble methods.
+The first part of the experiments evaluates the performance of each model under different data representations, as shown in Table 2.
+The results indicate that incorporating graph structures and service interactions generally enhances model performance, particularly in recall and F1 score.
+Table 2: Performance of each model across different data representations. The best performance is highlighted in bold.
 Model Dataset Phase F1 Score Precision Recall
@@
 D3 0.22018 0.13043 0.70588
-In order to effectively combine multiple models, the second part of the experiments evaluates the performance of different ensemble methods, as shown in Table 2.
-Different ensemble strategies are tested, such as combining XGB, SVM, stacked layers of the Self Attention model, and Isolation Forest.
-In the evaluation of the ensemble methods, we have compared two different methods, soft voting, hard&unanimous voting and hard&majority voting.
-Table 2 shows the best performances obtained.
-Overall, the soft voting approach provided better results for our objectives.
-Table 2: Performance metrics for ensemble methods across different phases.The best performance are highlighted in bold.
+The second part of the experiments evaluates the performance of different ensemble methods, as shown in table 3.
+We tested combinations of XGB, SVM, Self-Attention, and Isolation Forest, along with different voting mechanisms: soft, hard-unanimous, and hard-majority.
+Table 3 shows that the best performance is achieved with the combination of SVM, XGB and SA in hard-unanimous voting.
+Table 3: Performance of different ensemble methods.The best performance is highlighted in bold.
 Ensemble Model Method Dataset Phase F1 Score Precision Recall
@@
 XGB + SVM + SA + IF hard,unanimous D2+D3 0.36364 0.2963 0.47059
-XGB + SVM + SA soft D2+D3 0.51429 0.5 0.52941
-The individual models using the three representations mentioned in Section 3 and the results using ensemble learning both significantly outperforms the baseline result  [^3] The baseline result is from https://eval.ai/web/challenges/challenge-page/2267/overview, demonstrating the potential of our approach.
+XGB + SVM + SA hard,unanimous D2+D3 0.51429 0.5 0.52941
+Figure 4: The figure shows the F1 scores of the models across the three dataset levels, as well as the F1 scores obtained after applying ensemble learning methods.
+The individual models include XGB, SVM, TCN, LSTM, GRU, and the Self-Attention Model.
+The models used in the ensemble learning approach are XGB, Isolation Forest, SVM, and the Self-Attention Model.
+Fig.
+4 shows that the one-hop representation, two-hop representation as well as the ensemble learning methods all significantly outperforms the baseline result  [^5] The baseline result is from https://eval.ai/web/challenges/challenge-page/2267/overview, demonstrating the potential of our approach.
 5 Conclusion
-In this paper, we investigated the effectiveness of different machine learning and deep learning models on the challenging task of anomaly detection on dynamic knowledge graphs, with a specific focus on a microservices environment.
-Using a structured experimental approach, we analysed the performance of the models in three different representations: sequential data modelling, one-hop graph structure, and two-hop graph structure.
-Our results show that incorporating graph structures significantly improves the performance of anomaly detection models.
-In particular, the Two-Hop Graph Structure, which includes service interactions, produced the best results across most metrics, highlighting the importance of capturing complex interdependencies within the data.
-Deep learning models, particularly the Two-Stacked Self-Attention (2StackedSA) model, showed superior performance in leveraging these structures.
-Furthermore, the ensemble methods provided a substantial boost in performance by combining the strengths of top-performing individual models.
-Among the ensemble strategies, the Unanimous Voting approach, which requires all models to agree on the prediction of an anomaly, outperformed the Majority Voting approach in terms of precision.
-This aligns well with the competition’s emphasis on precision, highlighting the effectiveness of this conservative yet accurate method.
-Future work should address several areas for improvement.
-Firstly, improving the scalability of our models to handle larger and more complex datasets would be crucial.
-Secondly, exploring more advanced ensemble techniques, such as stacking or blending, could further improve detection accuracy.
-Thirdly, the integration of real-time anomaly detection capabilities would be highly beneficial for practical applications in dynamic environments.
-Finally, our study was limited to binary classification (normal/abnormal).
-Future research should focus on improving the accuracy of predicting different types of anomalies.
-Furthermore, although we used graph structure information, we did not apply Graph Neural Networks (GNNs), which could be a promising direction for future investigation and testing.
-This study demonstrates the significant potential of combining graph-based data representations with advanced machine learning and deep learning techniques for anomaly detection on dynamic knowledge graphs.
-By refining these methods and exploring the future directions outlined, we can develop more robust, accurate and scalable solutions for complex anomaly detection tasks in various domains.
+In this paper, we explored the effectiveness of various machine learning and deep learning models for anomaly detection on dynamic knowledge graphs, focusing on a microservices environment.
+Using a structured experimental approach, we analysed the performance of the models on three representations: sequential data, one-hop, and two-hop graph structures.
+Our results show that incorporating graph structures significantly enhances model performance.
+The two-hop graph structure, which includes service interactions, delivered the best results across most metrics, emphasising the importance of capturing complex interdependencies.
+Deep learning models, especially the Self-Attention model, performed particularly well in leveraging these structures.
+Furthermore, the ensemble methods significantly boosted performance by combining the strengths of the top individual models.
+Unanimous voting, which requires all models to agree on an anomaly, outperformed majority voting in terms of precision, matching the competition’s emphasis on precision.
+Future work should focus on several improvements.
+First, improving model scalability to handle larger, more complex datasets is essential.
+Second, exploring advanced ensemble techniques such as stacking or blending could improve detection accuracy.
+Third, integrating real-time anomaly detection would benefit practical applications in dynamic environments.
+Additionally, our study was limited to binary classification; future research should improve the accuracy of detecting different types of anomalies.
+Finally, although we used graph structure information, the application of graph neural networks remains a promising direction for further investigation.
+This study highlights the potential of combining graph-based data representations with advanced machine learning and deep learning techniques for anomaly detection on dynamic knowledge graphs.
+Refining these methods and exploring the suggested future directions can lead to more robust, accurate, and scalable solutions for complex anomaly detection in various domains.

### v3

@@
 Abstract
-In this paper, we explore different approaches to anomaly detection on dynamic knowledge graphs, specifically in a microservices environment for Kubernetes applications.
-Our approach explores three dynamic knowledge graph representations: sequential data, one-hop graph structure, and two-hop graph structure, with each representation incorporating increasingly complex structural information.
-Each phase includes different machine learning and deep learning models.
+In this paper, we explore different approaches to anomaly detection on dynamic knowledge graphs, specifically in a Micro-services environment for Kubernetes applications.
+Our approach explores three dynamic knowledge graph representations: sequential data, hierarchical data and inter-service dependency data, with each representation incorporating increasingly complex structural information of dynamic knowledge graph.
+Different machine learning and deep learning models are tested on these representations.
 We empirically analyse their performance and propose an approach based on ensemble learning of these models.
@@
 Dynamic Knowledge Graphs (DKG), as an intersection of heterogeneous and dynamic graphs, have attracted increasing attention due to their ability to represent the evolving nature of knowledge over time.
-By exploiting the ability to capture topological and attribute transformations, DKGs provide valuable insights into the evolution of knowledge in diverse domains such as social media [23], the Internet of Things (IoT) [22, 21], and e-commerce [^2] https://eval.ai/web/challenges/challenge-page/2267/overview.
-On the leading Micro-Services Architecture (MSA) platform, such as Kubernetes, there are different types of entities such as compute nodes, deployable units (pods), and services that work together to build applications such as e-commerce platforms.
-These data are typically stored in DKGs in the TTL format [^3] https://en.wikipedia.org/wiki/Turtle_(syntax).
-Predictive analysis based on such data is critical to the operation and optimisation of these platforms, particularly for anomaly detection tasks aimed at accurately identifying various types of anomalies.
-Despite rapid progress in commercial applications, as a relatively young field, there is still a lack of academic datasets and methodologies.
-This scarcity can be attributed to several challenges: anomaly events are typically very rare, resulting in an unbalanced data distribution that makes it difficult for models to effectively learn anomaly features and affects their generalisation ability in detecting new anomalies.
-In addition, models need to capture temporal and structural information and handle heterogeneous nodes/edges, which makes common RNNs and homogeneous GNNs inadequate.
-To address these challenges, this paper introduces a novel anomaly detection framework for dynamic knowledge graphs, leveraging both graph structure and temporal dependencies.
+By exploiting the ability to capture topological and attribute transformations, DKGs provide valuable insights into the evolution of knowledge in diverse domains such as Social Media [22], the Internet of Things (IoT) [21, 20], and E-commerce.
+Kubernetes-based Micro-services Architecture (MSA) platforms consist of interconnected entities, such as compute nodes, deployable units (pods), and services, which work together to build applications like e-commerce platforms.
+Since these entities are different and they also depend on each other through different relationships, the best way to represent the data information and the structural information is to construct a Knowledge Graph (KG) [40].
+Moreover, the information and dependencies among entities can change over time in MSA platforms.
+Therefore, the use of a Dynamic Knowledge Graph (DKG) [39] that records a complete snapshot of DKG at each time step is the best way to register and track these changes.
+In order to collect and store this DKG dataset, the organisers used the TTL format  [^2] https://en.wikipedia.org/wiki/Turtle_(syntax), a textual syntax for the Resource Description Framework (RDF) that allows KG information to be represented in a compact textual form, such as a triple consisting of a subject, a predicate and an object  [^3] https://www.w3.org/TR/rdf12-turtle/#language-features.
+Using the Turtle language, the entities and relations of a snapshot of each time step are stored in ternary relations.
+Predictions based on this structured data is crucial for the operation and optimisation of MSA.
+This ternary structure representation is similar to the syntax of natural language, which helps to improve the efficiency of querying and parsing, thus improving the effectiveness and robustness of the anomaly detection model.
+Although DKGs are well suited for modelling different nodes, variant relationships, and temporal information in Kubernetes-based MSA platforms, their analysing and learning approaches are not well explored yet.
+In addition, several challenges are specific to the anomaly detection task of Kubernetes-based MSA platforms: 1.
+Anomalous events are usually very rare, resulting in an imbalanced data distribution, which prevents the model’s ability to learn and generalise anomalous features effectively.
+2.
+Traditional RNNs and homogeneous GNNs are insufficient for accurate anomaly detection due to the need to capture temporal and structural information in Kubernetes environments, coupled with the presence of heterogeneous nodes and edges.
+To address these challenges, we need to develop specialised approaches capable of handling the dynamic and complex structures of Kubernetes-based DKGs.
+This paper introduces a novel anomaly detection framework for dynamic knowledge graphs, leveraging both graph structure and temporal dependencies.
 Our method integrates neighbourhood feature extraction with advanced machine learning techniques, including Support Vector Machine (SVM) [6], eXtreme Gradient Boosting (XGB) [8], and Self-Attention mechanism (SA) [13] to enhance the model’s ability to capture complex patterns of anomalies.
@@
 The structure of the paper is as follows.
-Section 2 gives the formulation of the problem and describes related work.
+Section 2 gives the problem formulation and describes related work.
 Section 3 presents the methodology of the paper.
 Section 4 describes the experimental part of the work and discusses the results obtained.
-Section 5 concludes our contribution to the ADDKG challenge and discusses future work.
+Section 5 concludes our contribution to the ADDKG challenge and discusses future works.
 2 Related Work
@@
 Among them, the edge list is widely used in software development as it requires less memory.
-An edge list is a collection of edges, each represented as a pair of nodes.
-In the realm of dynamic graphs, existing research categorises them into Continuous-Time Dynamic Graphs (CTDG) and Discrete-Time Dynamic Graphs (DTDG).
+An edge list is a collection of edges.
+Each edge is represented as a pair of nodes.
+In the realm of dynamic graphs, existing research categorises them into Continuous-Time Dynamic Graphs (CTDG) and Discrete-Time Dynamic Graphs (DTDG) [5].
 CTDGs represent dynamic graphs using timestamped events, such as $(u,v,t)$, where $u$ and $v$ are nodes and $t$ is the timestamp of the interaction.
-In contrast, DTDGs represent dynamic graphs as a sequence of static graph snapshots $\{G_{1},G_{2},...,G_{T}\}$ at different time intervals.
+In contrast, DTDGs represent dynamic graphs as a sequence of static graph snapshots $\{G_{1},G_{2},...,G_{T}\}$.
 Combined with the dynamic graph representations, the Dynamic Knowledge Graph (DKG) is modelled as a heterogeneous dynamic graph, represented by timestamped edge lists.
-Among the various formats for representing DKGs, Turtle (TTL) is one of the most commonly used.
-TTL allows the serialisation of nodes, edges, their attributes and timestamps, making it a versatile format for representing entities and relationships in a multi-agent system (MSA).
-An example of a TTL-based DKG is given in Eq.
-1, where $s$ is the subject node, $o$ is the object node, $p$ is the predicate (or relationship), and $t$ is the timestamp.
-$\text{DKG}=\{(s,p,o,t)\mid t\in[t_{1},t_{2},...,T]\}$ (1)
-In the problem studied in this paper, anomalous data is represented by its entity $s$, the time interval during which the anomaly occurs $(t_{start},t_{end})$, and the class of the anomaly $c$.
-Table 1 shows all the anomaly types $c$ in the dataset.
+Among the various formats for representing DKGs, Turtle format (TTL) is used for this competition.
+The TTL format allows a snapshot of DKGs to be modelled as a series of triples in the structure of (subject, predicate, object).
+Then a filename that records a timestamp is bound to this snapshot for registering the temporal information.
+An example of a TTL-based DKG at moment $t$ is given in Eq.
+2.1, where $S$ is the ensemble of subject nodes, $O$ is the ensemble of object nodes, $P$ is the ensemble of predicates (or relationships), and $T$ is the ensemble of timestamps.
+$\text{DKG}_{t}=\{(s,p,o)\mid s\in S,p\in P,o\in O\}$
+Timestamp $\text{Timestamp }t\text{ is associated with DKG}_{t},t\in T$ (1)
+In the problem studied in this paper, all anomalies are labelled at the graph level. Specifically, each anomaly in the DKG is labelled with a category $c$ occurring from $t_{start}$ to $t_{end}$.
 Table 1: Description of system faults and their types.
@@
 However, these methods often struggle with high-dimensional data and complex dependencies [26].
-To overcome these problems, common machine learning approaches include Support Vector Machine (SVM) [6], eXtreme Gradient Boosting (XGB) [8], and Isolation Forest (IF) [7].
+To overcome these problems, common machine learning approaches include SVM [6], XGB [8], and Isolation Forest (IF) [7].
 SVM identifies the optimal hyperplane that maximises the distance between normal data points and the decision boundary, effectively distinguishing outliers.
 XGB is a gradient-boosting decision tree algorithm that is able to evaluate the importance of each feature and find those that are most effective in separating normal from abnormal data.
-Isolation Forest is an unsupervised method that detects anomalies by constructing decision trees to isolate observations, with outliers requiring fewer splits to isolate due to their rarity and distinctiveness.
+IF is an unsupervised method that detects anomalies by constructing decision trees to isolate observations, with outliers requiring fewer splits to isolate due to their rarity and distinctiveness.
 With the advancement of software algorithms and hardware, deep learning methods have been increasingly popular for anomaly detection due to their ability to extract complex features [15].
@@
 TCNs handle long-term dependencies through causal convolutions, while SA captures long-range dependencies by examining relationships across sequences.
-For graph anomaly detection, Graph Neural Networks (GNN) propagate information through graph structures to identify anomalous patterns [19, 18].
-Despite advancements in anomaly detection algorithms, significant challenges persist in dynamic knowledge graphs.
-Traditional machine learning methods struggle to capture the complex temporal and structural dependencies, while deep learning models require large amounts of labelled data, which is often scarce.
-Although techniques like negative sampling [19, 18, 20] can generate anomalous edges in dynamic graphs without attributed edges, this approach is infeasible for dynamic knowledge graphs with attributed edges.
+With the development and popularity of financial, social, traffic and biological networks, graph-based models have gained significant attention in anomaly detection tasks.
+Graph-based anomaly detection can be classified into four types [27]: node anomalies, edge anomalies, subgraph anomalies and full graph anomalies.
+Early research [28, 29, 30, 31] focused on static graph anomaly detection, mainly on node anomalies, with limited attention to edge and subgraph anomalies.
+However, real-world graph data tends to change over time, which introduces additional complexity.
+To address this problem, researchers have proposed dynamic graph solutions that aim to detect anomalies that occur as the graph changes.
+Some studies [17, 18, 19] have explored how to detect anomalies at edges or nodes in dynamic graphs.
+Moreover, networks not only evolve over time, but also present complex structures with heterogeneous nodes and edges.
+Knowledge Graphs (KGs) provide a powerful framework for modelling such heterogeneous graphs, making them suitable for anomaly detection tasks.
+Despite the potential of KGs, the application of anomaly detection to (dynamic) KGs is still limited, mainly due to the unstructured nature of KGs and the inherent challenges of dealing with their dynamic and heterogeneous nature.
+The growth of cloud services and distributed systems has led to the widespread use of MSA in software and web applications, moving away from traditional monolithic architectures [32].
+MSA breaks down an application into independent services, each responsible for a specific function and able to communicate.
+This architecture enhances the scalability and flexibility of software, but as the number of services increases, it becomes increasingly difficult to manually identify anomalies.
+Therefore, an automated anomaly detection system is essential to maintain the system.
+Most existing work in this area leverages three main data sources (logs [33], performance metrics [34], and distributed traces [35]) and uses machine learning algorithms such as SVMs, K-Nearest Neighbours (KNN) [36], Random Forests (RF) [37], and Multi-Layer Perceptrons (MLPs)[38] for anomaly detection.
+However, these methods often ignore the structural relationships between services.
+Modelling MSA as (dynamic) KGs can capture these inter-service dependencies and provide a more comprehensive framework for anomaly detection.
+However, the integration of DKGs with MSA for anomaly detection is still under-researched, mainly due to the challenges posed by the complexity and scalability of DKGs.
+In addition to all these, the anomaly detection task itself presents challenges.
+Traditional machine learning methods struggle to capture the complex temporal and structural dependencies, while deep learning models require a large amount of labelled data and a balance between the amount of normal and abnormal data.
+However, in practice, the amount of anomalous data tends to be much less than the amount of normal data [27].
+Although techniques like negative sampling [18, 17, 19] can generate anomalous edges in dynamic graphs without attributed edges, this approach is infeasible for DKGs with attributed edges.
 Capturing both temporal and structural dependencies remains key to tackling this challenge.
+To address these gaps, the ADDKG competition provides a dataset based on MSA and DKGs and encourages participants to develop innovative solutions for anomaly detection in this complex environment.
 3 Methodology
-This section introduces three representations of dynamic knowledge graphs: sequential data, one-hop, and two-hop graph structures, each adding increasing structural complexity.
+This section introduces three representations of dynamic knowledge graphs: sequential data, hierarchical data, and inter-service dependency data, each adding increasing structural complexity.
 We then describe the machine learning and deep learning methods applied to these representations, followed by proposed ensemble learning schemes.
-Figures 1 and 2 provide an overview of the pipeline, including the data structure, the three levels of the data processing, and the final output.
+3.1 Simplified Tree Structure
 Figure 1: Left: The ontology overview of the relationships between Kubernetes concepts of the ADDKG dataset.
@@
 The Connections serve as attributes of the edges between Pods and Services.
-Figure 2: This figure presents the methodology used in this study, with three levels of processing: sequential data, one-hop, and two-hop graph structures.
-Each level introduces increasing structural complexity, showing how these representations are fed into anomaly detection models.
-3.1 DKG as Sequential Data
-The time-evolving attributes of nodes can be obtained by extracting raw TTL data, without considering the structural information of the DKG.
-As shown in the "DKG as Sequential Data" part of Fig.
-2, we merge all the sequential data of all the vertices in a DKG by the timestamp.
-This includes vertices of type $\mathcal{C}$ representing "Cluster", type $\mathcal{P}$ representing "Pod", type $\mathcal{S}$ representing "Service", and type $\mathcal{N}$ representing "Node".
-For each given node category $k$, the time-evolving attributes can be represented as a tensor $\mathbf{\tilde{X}}_{k}\in\mathbb{R}^{N_{k}\times T\times d_{k}}$, where $N_{k}$ indicated the number of instances of the category $k$, $T$ indicates the number of snapshots, and $d_{k}$ indicates the number of features.
-A sliding window then captures temporal dependencies by concatenating the historical information of the past $\tau$ time steps, as shown in Eq.
+The left part of Fig.
+1 illustrates the ontology of the DKG utilised in the ADDKG dataset.
+The ontology comprises a total of ten distinct vertices (also known as functional units in ADDKG DKGs): Cluster, Node, Pod, Container, Image, Connection, Service, Workload, ReplicaSet and Deployment.
+These vertices are connected to each other with different relationships.
+Through a comprehensive analysis, three principal observations were identified:
+Primary Vertices.
+The ontology includes three primary vertices—Node, Container, and Connection.
+These are designated as primary vertices because only their properties record CPU performance metrics: minimum, maximum, mean, deviation, variance, 50th percentile, 90th percentile, and 95th percentile of CPU resource consumptions at each time step.
+Directional Relationships.
+All edges within the ontology are unidirectional.
+For example, relationships such as (Cluster, hasNode, Node) and (Node, hasPod, Pod) demonstrate this unidirectional nature.
+The only exception is the bidirectional edge pair "controls" and "is controlled by," which essentially represent the same relationship from opposite perspectives.
+Simplified Tree Structure.
+The DKG can be abstracted into a simplified tree structure, as depicted in the right part of Fig.
+1.
+This simplification consolidates information into three main vertices: Node, Pod, and Service, connected by the relationships (Node, hasPod, Pod) and (Pod, hasConnection, Service).
+According to the ontology, node attributes from the primary vertices (Container and Connection) can be effectively propagated to Pod and Service through the "hasContainer" and "hasConnection" relations.
+The illustrated tree structure in Fig.
+1, comprising Node, Pod, and Service, serves as a reduced version of the tree structure, hence it can not represent the whole DKG.
+We will use this example to facilitate the later explanation of the three data representation methods used in our study.
+In the example, yellow vertices represent Nodes, pink vertices represent Pods, and green vertices represent Services.
+Each vertex type can have varying numbers of instances; for instance, this example includes one Node (n1), three Pods (p1, p2, p3), and four Services (s1, s2, s3, s4).
+3.2 Sequential Data Representation
+In this phase, the representation of the DKGs can be constructed by concatenating the time-evolving attributes of the vertices in the simplified tree structure, specifically from vertices Node, Pod and Service.
+Figure 2: This figure presents the sequential data representation based on the simplified tree structure.
+Operationally, time-evolving attributes can be extracted by querying the nodes and relations from the triples in TTL files.
+We can then simply merge these data by timestamps to form the sequential data representation.
+This operation largely ignores the structural relations between nodes, e.g., (Pod, hasNode, Node).
+As shown in the "Sequential Data Representation" part of Fig.
+2, temporal data from Node 1 is systematically merged into Pods 1, 2, and 3 based on their respective timestamps.
+Then the merged data from these Pods is further consolidated into Services 1, 2, 3, and 4, ensuring temporal consistency throughout the hierarchical aggregation process.
+Mathematically, for each given vertex category $k$, the time-evolving attributes can be represented as a tensor $\mathbf{\tilde{X}}_{k}\in\mathbb{R}^{T\times D_{k}}$, where $T$ indicates the number of snapshots, and $D_{k}$ indicates the number of features of category $k$ by concatenating all the instances, which equals to $N_{k}\times d_{k}$ with $N_{k}$ indicates the number of instances of the category $k$, and $d_{k}$ is the number of features of the category $k$.
+Then we can merge all vertices Node, Pod and Service together based on timestamps, thus we can get a $\mathbf{\tilde{X}}\in\mathbb{R}^{T\times D}$, where $D=D_{Node}+D_{Pod}+D_{Service}=\sum^{3}_{k=1}N_{k}*d_{k}$.
+A sliding window then captures temporal dependencies by concatenating the historical information of the past H time steps, as shown in Eq.
 2.
-$\mathbf{f}(t_{j})=[\mathbf{x}(t_{j}-\tau+1)\,|\,\mathbf{x}(t_{j}-\tau+2)\,|\,\cdots\,|\,\mathbf{x}(t_{j})]$ (2)
-Therefore, the sequential data representation of a given category $k$ in the DKG can be expressed by a matrix $\mathbf{X}\in\mathbb{R}^{T\times D_{k}}$, where $D_{k}=N_{k}\times d_{k}$, as shown in Eq.
+$\mathbf{f}(t_{j})=[\mathbf{\tilde{x}}(t_{j}-H+1)\,|\,\mathbf{\tilde{x}}(t_{j}-H+2)\,|\,\cdots\,|\,\mathbf{\tilde{x}}(t_{j})]$ (2)
+Therefore, the sequential data representation the DKG can be expressed by a matrix $\mathbf{X}\in\mathbb{R}^{T\times D\times H}$, where $H$ is the number of historical steps, as shown in Eq.
 3, where $\mathbf{f}(t_{j})$ is the aggregated attribute vector at timestamp $t_{j}$.
-$\mathbf{X}=\begin{bmatrix}\mathbf{f}(t_{1})\\ \mathbf{f}(t_{2})\\ \vdots\\ \mathbf{f}(t_{T})\end{bmatrix}=\begin{bmatrix}\mathbf{x}(t_{1}-\tau+1)&\mathbf{x}(t_{1}-\tau+2)&\cdots&\mathbf{x}(t_{1})\\ \mathbf{x}(t_{2}-\tau+1)&\mathbf{x}(t_{2}-\tau+2)&\cdots&\mathbf{x}(t_{2})\\ \vdots&\vdots&\ddots&\vdots\\ \mathbf{x}(t_{T}-\tau+1)&\mathbf{x}(t_{m}-\tau+2)&\cdots&\mathbf{x}(t_{m})\end{bmatrix}$ (3)
-Once the matrix representation of the attributes $\mathbf{X}$ is obtained, additional information can be computed.
-For example, the difference captures the trend of the time series, while the variance reflects the stability of the data fluctuations.
-Based on this data representation, the temporal information across multiple time steps is aggregated into a single vector.
-Consequently, machine learning algorithms such as SVM, XGB, and IF, as well as deep learning models like MLP, TCN, and RNN, can use this input to capture information across multiple time steps.
-3.2 One-Hop Graph Structure
-To explore the structural information of the knowledge graph, we propose to use one-hop neighbour aggregation to capture the relationships between different entities.
-Given that the relationships in the DKG of an MSA are mainly hierarchical, i.e., a cluster contains multiple nodes, a node contains multiple pods, and a pod is connected to services through multiple connections (see Fig.
-1), it is possible to aggregate information from leaf nodes to root nodes to obtain the structural-temporal features of each node at each timestamp, as shown the aggregate method in the "One-Hop Chart Structure" section of Fig.
-2.
-$S(t)=\bigcup_{i=1}^{n_{S}}\left(S_{i}(t),\bigcup_{j\in C(S_{i})}\left(C_{j}(t),\bigcup_{k\in P(C_{j})}\left(P_{k}(t),\bigcup_{l\in N(P_{k})}N_{l}(t)\right)\right)\right)$ (4)
-For instance, to encode and predict data for the category Service, information related to the Service can be aggregated.
-Specifically, for all service instances $\mathcal{S}(t)$ at timestamp $t$, the dynamic information from related Nodes, Pods, Containers and Connections can be aggregated to form the attributes of the Service, as shown in Eq.
-4, where $n_{S}$ represents the number of Service instances, $S_{i}$ represents the $i$-th instance of the Service category, $C_{j}(S_{i})$ represents the elements in the set of Connection instances related to $S_{i}$, $P_{k}(C_{j})$ represents the elements in the set of Pod instances related to $C_{j}$, and $N_{l}(P_{k})$ represents the elements in the set of Node instances related to $P_{k}$.
-In this manner, multiple one-hop graph structures are used for message propagation, and these structural-temporal features provide the model with a richer context.
-3.3 Two-Hop Graph Structure
-Some graph machine learning algorithms have pointed out that a larger convolutional field, or aggregation of multi-hop neighbours, can optimise the model.
-Inspired by the potential of aggregating second-hop neighbours to provide richer information, we observed that in the tree-like graph of the MSA problem, considering only the current level and its root forms a bipartite graph.
-For nodes of category $k$, the two-hop neighbours remain nodes of the same category.
-Recognising potential interdependencies between Service instances at each timestamp, we propose two-hop neighbour aggregation.
-This can be intuitively achieved by concatenating vertex features, as shown in Eq.
-5 where $N_{\mathcal{S}}$ represents the number of Service instances and $S_{i}(t)$ represents the data for the $i$-th Service instance at timestamp $t$.
+$\mathbf{X}=\begin{bmatrix}\mathbf{f}(t_{1})\\ \mathbf{f}(t_{2})\\ \vdots\\ \mathbf{f}(t_{T})\end{bmatrix}=\begin{bmatrix}\mathbf{\tilde{x}}(t_{1}-H+1)&\mathbf{\tilde{x}}(t_{1}-H+2)&\cdots&\mathbf{\tilde{x}}(t_{1})\\ \mathbf{\tilde{x}}(t_{2}-H+1)&\mathbf{\tilde{x}}(t_{2}-H+2)&\cdots&\mathbf{\tilde{x}}(t_{2})\\ \vdots&\vdots&\ddots&\vdots\\ \mathbf{\tilde{x}}(t_{m}-H+1)&\mathbf{\tilde{x}}(t_{m}-H+2)&\cdots&\mathbf{\tilde{x}}(t_{m})\end{bmatrix}$ (3)
+Once the matrix representation of the attributes $\mathbf{X}$ is obtained, additional time-dependent features can be computed to enhance the model’s ability to detect anomalies.
+For example, calculating the difference between consecutive data points can capture the trend of the time series and understand the direction of change.
+Similarly calculating the variance of data within a specific time window can reflect the stability of the data.
+By additionally incorporating the difference and variance as temporal features, the dynamic behaviour of the system can be effectively captured and the ability to identify anomalies can be improved.
+Based on this data representation, machine learning algorithms such as SVM, XGB, and IF, as well as deep learning models like MLP, TCN, LSTM, GRU and SA, can use this input to capture information across multiple time steps.
+3.3 Hierarchical Data Representation
+Figure 3: This figure presents the hierarchical data representation based on the simplified tree structure.
+To investigate the structural information within the DKGs, we propose a hierarchical neighbourhood information aggregation method designed to capture the connectivity relationships among various entities.
+In the simplified tree structure, hierarchical relationships exist between vertices.
+For example, a Node has multiple Pods and a Pod is connected to different Services through multiple Connections and the "hasConnection" relationship (see Fig.
+1).
+Operationally, when relationships exist across Nodes, Pods, and Services, we can progressively aggregate their instances’ time-evolving attributes from Nodes to Services using aggregation methods such as mean, max, and min.
+This approach enables the construction of a structural-temporal representation for each Service instance at each timestamp, as illustrated in the "Hierarchical Data Structure" in Fig.
+3.
+This operation influences the feature dimension $D$ of $\mathbf{X}$, defined in Section 3.2.
+Initially, $D$ was determined by $N_{k}$, the number of instances of category $k$.
+However, we have now redefined $D$ into $\sum^{3}_{k=1}AGG*d_{k}$, with $AGG$ corresponding to the number of aggregation methods employed.
+This aggregation process is analogous to a multi-hop operation, where information is aggregated from multiple hops away from the target Service.
+For instance, for Service 1 Fig.
+3, temporal data is first aggregated from Node 1, followed by the aggregation of data from Pods 1, 2, and 3, and finally merging them based on the timestamp.
+This method ensures that the temporal dynamics and structural relationships inherent in the DKG are effectively captured, thereby enhancing the accuracy and robustness of anomaly detection within the MSA.
+3.4 Hierarchical and Inter-Service Representation
+Some graph machine learning algorithms have pointed out that a larger convolutional field can optimise the model [5].
+Since we have already aggregated the multi-hop structural-temporal information into the Service vertex, in this section we supposed that all Service instances can influence each other.
+Therefore, we can further enrich the DKG representations by providing additional correlations between Service instances.
+Figure 4: This figure presents the inter-service data representation based on the simplified tree structure.
+This representation can be intuitively achieved by concatenating vertex features, as shown in Eq.
+4 where $N_{\mathcal{S}}$ represents the number of Service instances and $\mathbf{s}_{i}(t)$ represents the data for the $i$-th Service instance at timestamp $t$.
+However, this approach will greatly increase the size of the feature dimension $D$ of $\mathbf{X}$, requiring more computational resources during feature processing and model training.
 $\mathbf{X}_{\text{service}}(t)=[\mathbf{s}_{1}(t),\mathbf{s}_{2}(t),\ldots,\mathbf{s_{\mathit{N_{S}}}}(t)];$ (4)
-Another method of capturing two-hop neighbours is using the Self-Attention mechanism (SA) to capture the correlations among vertices.
+Another method is adding implicitly the inter-service dependencies using the model layers, like Self-Attention, LSTM to capture the correlations among vertices during the model training.
 SA works by computing the relevance (or attention) between each pair of elements in a sequence, enabling the model to weigh the importance of each element relative to others.
 This allows the model to focus on relevant parts of the input sequence when making predictions, effectively capturing long-range dependencies and complex relationships within the data.
-See the "two hop graph structure" part in Fig.
-2.
+See the "Inter-Service Dependencies" part in Fig.
+4.
 3.5 Ensemble Learning
@@
 It allows multiple models to make predictions on the same problem and vote on the final predictions.
-Voting can be hard or soft voting, with the mechanism such as unanimous and majority voting, which will be introduced in this subsection.
+Voting can be hard or soft voting, with mechanisms such as unanimous and majority voting, which will be introduced in this subsection.
 Figure 5: The ensemble learning framework used in this competition.
-The best-performing models: XGBoost (XGB), Support Vector Machine (SVM), and Self-Attention (SA) are combined using three voting mechanisms: hard-unanimous, hard-majority, and soft voting, to generate the prediction $\hat{y_{1}}$, where $\hat{y_{1}}\in\{0,1\}$, with 0 indicating normal and 1 indicating anomaly.
+The best-performing models: XGB, SVM, and SA are combined using three voting mechanisms: hard-unanimous, hard-majority, and soft voting, to generate the prediction $\hat{y_{1}}$, where $\hat{y_{1}}\in\{0,1\}$, with 0 indicating normal and 1 indicating anomaly.
 Optionally, a two-stage model combination can be applied, where the unsupervised Isolation Forest further refines the anomalies detected in $\hat{y_{1}}$ to reduce false positives.
@@
 Both methods are tested on multiple sets of models and it are discussed in the next section.
+In the pipeline, we also explore an optional two-stage model to improve anomaly detection accuracy.
+In this approach, we first identify potential anomalies by merging the results of the three models through hard or soft voting, and then use these initial predicted anomalies as a new dataset $\mathbf{X}_{2}$, and subsequently feed $\mathbf{X}_{2}$ into the unsupervised algorithm Isolation Forest as a secondary filtering mechanism.
+The aim of incorporating IF into this two-stage pipeline is to reduce the number of false positives through tree-based segmentation and isolation of outliers.
+This approach allows us to identify unique, isolated anomalies that may have been inaccurately classified in the first stage, thereby improving the robustness of the overall anomaly detection process.
 4 Experiment
@@
 4.2 Performance of Models
-This section presents the experimental results for the three data representations: D1, D2, and D3, correspond respectively to sequential data, one-hop, and two-hop graph representations.
+This section presents the experimental results for the three data representations: D1, D2, and D3, correspond respectively to sequential data, hierarchical data, and inter-service data.
 Table 2 summarises the performance of individual models, while table 3 focuses on the results of different ensemble methods.
 The first part of the experiments evaluates the performance of each model under different data representations, as shown in Table 2.
-The results indicate that incorporating graph structures and service interactions generally enhances model performance, particularly in recall and F1 score.
+Table 2 only showcases the best results for each case.
+It is worth remarking that the baseline score provided by the challenge organisers is the score obtained by predicting all snapshots given at each timestamp as anomalies, and thus it has a recall equal to 1.
+The results indicate that incorporating graph structures and service interactions generally enhances model performance, particularly in recall and F1 scores.
+While our study aimed to evaluate multiple models comprehensively across different configurations, certain constraints affected the depth of testing.
+Due to daily submission limits, some models could not be exhaustively tested across all configurations.
+For instance, XGB and SVM were only evaluated under configuration D2.
+D3 representation needs to incorporate additional inter-service information, which requires significant computational resources, limiting our ability to test them further.
+Moreover, the TCN model detected no anomalies following training under D3; as a result, its findings are omitted from this analysis.
 Table 2: Performance of each model across different data representations. The best performance is highlighted in bold.
@@
 The second part of the experiments evaluates the performance of different ensemble methods, as shown in table 3.
-We tested combinations of XGB, SVM, Self-Attention, and Isolation Forest, along with different voting mechanisms: soft, hard-unanimous, and hard-majority.
-Table 3 shows that the best performance is achieved with the combination of SVM, XGB and SA in hard-unanimous voting.
+Table 3 only shows the best results for each case.
+We tested combinations of XGB, SVM, SA, and IF, along with different voting mechanisms: soft, hard-unanimous, and hard-majority.
+The results show that the combination of SVM, XGB and SA achieves the best performance in hard-unanimous voting.
 Table 3: Performance of different ensemble methods.The best performance is highlighted in bold.
@@
 XGB + SVM + SA hard,unanimous D2+D3 0.51429 0.5 0.52941
-Figure 4: The figure shows the F1 scores of the models across the three dataset levels, as well as the F1 scores obtained after applying ensemble learning methods.
-The individual models include XGB, SVM, TCN, LSTM, GRU, and the Self-Attention Model.
-The models used in the ensemble learning approach are XGB, Isolation Forest, SVM, and the Self-Attention Model.
+Figure 6: The figure shows the F1 scores of the models across the three data representations in the left, as well as the F1 scores obtained after applying ensemble learning methods in the right.
+The individual models include XGB, SVM, TCN, LSTM, GRU, and SA Model.
+The models used in the ensemble learning approach are XGB, IF, SVM, and SA Model.
 Fig.
-4 shows that the one-hop representation, two-hop representation as well as the ensemble learning methods all significantly outperforms the baseline result  [^5] The baseline result is from https://eval.ai/web/challenges/challenge-page/2267/overview, demonstrating the potential of our approach.
+6 shows that the hierarchical data representation, inter-service dependency representation as well as ensemble learning methods all significantly outperforms the baseline result, demonstrating the potential of our approach.
 5 Conclusion
-In this paper, we explored the effectiveness of various machine learning and deep learning models for anomaly detection on dynamic knowledge graphs, focusing on a microservices environment.
-Using a structured experimental approach, we analysed the performance of the models on three representations: sequential data, one-hop, and two-hop graph structures.
+In this paper, we explored the effectiveness of various machine learning and deep learning models for anomaly detection on dynamic knowledge graphs, focusing on a Micro-services environment.
+Using a structured experimental approach, we analysed the performance of the models on three representations: sequential data, hierarchical data, and inter-service dependency data.
 Our results show that incorporating graph structures significantly enhances model performance.
-The two-hop graph structure, which includes service interactions, delivered the best results across most metrics, emphasising the importance of capturing complex interdependencies.
-Deep learning models, especially the Self-Attention model, performed particularly well in leveraging these structures.
-Furthermore, the ensemble methods significantly boosted performance by combining the strengths of the top individual models.
+The inter-service structure, which includes service interactions and hierarchical representation, delivered the best results across most metrics, emphasising the importance of capturing complex structures and (inter-)dependencies.
+Deep learning models, especially the SA model, performed particularly well in leveraging these structures.
+Furthermore, the ensemble methods significantly improved performance by combining the strengths of the top individual models.
 Unanimous voting, which requires all models to agree on an anomaly, outperformed majority voting in terms of precision, matching the competition’s emphasis on precision.
@@
 First, improving model scalability to handle larger, more complex datasets is essential.
-Second, exploring advanced ensemble techniques such as stacking or blending could improve detection accuracy.
+Second, exploring advanced ensemble techniques such as stacking could improve detection accuracy.
 Third, integrating real-time anomaly detection would benefit practical applications in dynamic environments.
@@
 Finally, although we used graph structure information, the application of graph neural networks remains a promising direction for further investigation.
+Moreover, our data representations are high-dimensional, we can compare and use some low-dimensional representations, such as node embedding, graph embedding to test our proposed solutions.
 This study highlights the potential of combining graph-based data representations with advanced machine learning and deep learning techniques for anomaly detection on dynamic knowledge graphs.